Cassidy is 3 times as old as Drew. In 8 years, Cassidy will only be twice as old as Drew. How old is Cassidy right now?
Answer by alicealc(293)

Cassidy's age = C
Drew's age = D

Cassidy is 3 times as old as Drew:
C = 3D
In 8 years, Cassidy will only be twice as old as Drew
C + 8 = 2 * (D + 8)
3D + 8 = 2D + 16
3D - 2D = 16 - 8
D = 8
C = 3D = 3 * 8 = 24

so, Cassidy is 24 years old now